# Fluctuations for non-Hermitian dynamics

### v1

@@ 5.2.3 Wick Theorem
By (5.21) and the martingale representation theorem, we can write
$I_{2}(f^{(i)},0,t_{i})=\int_{0}^{t}\mathcal{C}_{s}^{1/2}{\mathrm{d}}\mathfrak{b}_{s}+\mathcal{O}(N^{-c}),$ (5.23)
for some small fixed $c>0$, where $\mathfrak{b}_{s}\in\mathbb{R}^{p}$ is a standard Brownian motion, independent of the $I_{1}(f^{(i)},0,t_{i})$’s, and $\mathcal{C}_{s}$ is a $p\times p$ matrix such that
@@
This shows that, modulo a negligible error $N^{-c}$, the $I_{2}(f^{(i)},0,t_{i})$ are jointly a multivariate Gaussian random variable with covariance $\int_{0}^{t}(\mathcal{C}_{s})_{ij}{\mathrm{d}}s$.
Now a Wick theorem as in (3.9) for the product (5.22) easily follows. In fact, (5.23) readily implies a Wick theorem for $I_{2}(f^{(i)},0,t_{i})$; we are thus left only with products of the $I_{1}(f^{(i)},0,t_{i})$’s . Then, by [CipErdSch2023, Proposition 3.3], we also conclude that the $I_{1}(f^{(i)},0,t_{i})$’s satisfy a Wick theorem as they consists of product of resolvents evaluated at the same time. This concludes the proof of of Proposition 3.3.
@@
[…]
@@
which together with 5.27 gives
@@ Appendix A Stieltjes transform
$|m(z)|\leqslant\frac{1}{{\rm dist}(z,\operatorname{supp}(\mu))},\quad\quad|m^{\prime}(z)|\leqslant\frac{{\rm Im}[m(z)]}{{\rm Im}[z]},\quad|m^{(p)}(z)|\leqslant\frac{p!{\rm Im}[m(z)]}{{\rm dist}(z,\operatorname{supp}(\mu))^{p-1}{\rm Im}[z]}.$
@@
$|m^{\prime}(z)|=\left|\int_{\mathbb{R}}\frac{{\rm d}\mu(x)}{(x-z)^{2}}\right|\leqslant\int_{\mathbb{R}}\frac{{\rm d}\mu(x)}{|z-x|^{2}}=\frac{1}{{\rm Im}[z]}\int_{\mathbb{R}}\frac{{\rm Im}[z]{\rm d}\mu(x)}{|z-x|^{2}}=\frac{{\rm Im}[m(z)]}{{\rm Im}[z]}.$
@@
$|m^{(p)}(z)|=\left|\int_{\mathbb{R}}\frac{p!{\rm d}\mu(x)}{(x-z)^{p+1}}\right|\leqslant\int_{\mathbb{R}}\frac{p!{\rm d}\mu(x)}{|z-x|^{p+1}}\leqslant\int_{\mathbb{R}}\frac{p!{\rm d}\mu(x)}{|z-x|^{2}{\rm dist}(z,\operatorname{supp}(\mu))^{p-1}}=\frac{p!{\rm Im}[m(z)]}{{\rm dist}(z,\operatorname{supp}(\mu))^{p-1}{\rm Im}[z]},$
@@
$|m^{\prime}(z)|\lesssim\max|\widetilde{m}^{\prime}(u)|+\frac{\sqrt{\phi{\rm Im}[m(z)]}}{\eta}.$ (A.2)
@@
$|\partial_{u}{\rm Im}[m(u)]|\leqslant|m^{\prime}(u)|\leqslant|\frac{{\rm Im}[m(u)]}{{\rm Im}[u]},$
@@
$|m(w)-m(z)-(w-z)m^{\prime}(z)|\leqslant\frac{1}{2}|w-z|^{2}\sup|m^{\prime\prime}(u)|$
@@
$|m^{\prime}(z)|$ $\leqslant\frac{|m(w)-m(z)|}{|w-z|}+\frac{1}{2}|w-z|\sup|m^{\prime\prime}(u)|\lesssim\max|\widetilde{m}^{\prime}(u)|+\frac{\phi}{|w-z|}+|w-z|\frac{{\rm Im}[m(z)]}{{\rm Im}[z]^{2}}$
$\lesssim\max|\widetilde{m}^{\prime}(u)|+\frac{\sqrt{\phi{\rm Im}[m(z)]}}{{\rm Im}[z]},$
@@ Appendix B Well-posedness
Still following [RogShi1993], consider $\phi({\bf{x}})=-\sum^{*}_{i,j}\log|x_{i}-x_{j}|$, where here and in the following $\sum_{i,j}^{*}$ (resp. $\sum^{*}_{i}$) means summation over couples $-N\leqslant i\neq j\leqslant N$ such that $i,j\neq 0$ (resp. $-N\leqslant i\leqslant N$ such that $i\neq 0$). Let ${\bf{y}}_{t}:={\bf{x}}_{t\wedge\tau_{\varepsilon}}$, then Itô’s formula gives
@@
so that the above parenthesis is smaller than ($\sum^{*}_{i,j,k}$ below means summation over all $-N\leqslant i,j,k\leqslant N$ such that $i,j,k$ are all distinct and non-zero)
$-\sum_{i}^{*}\left(\sum^{*}_{j:j\neq i}\frac{1}{y_{i}-y_{j}}\right)^{2}+\sum^{*}_{i,j}\frac{1}{(y_{i}-y_{j})^{2}}=-\sum^{*}_{i,j,k}\frac{1}{(y_{i}-y_{j})(y_{i}-y_{k})}=0,$

### v2

@@ 5.2.3 Wick Theorem
By the martingale representation theorem (see e.g. [Kal02, Theorem 18.12]), we can write (recall the definition of $I_{2}$ from below (5.3))
$I_{2}(f^{(i)},0,t_{i})=\int_{0}^{t_{i}}\big(\widehat{\mathcal{C}_{s}}^{1/2}{\mathrm{d}}\mathfrak{b}_{s}\big)_{i},\qquad\quad i\in[p],$ (5.23)
where $\mathfrak{b}_{s}\in\mathbb{R}^{p}$ is a standard real $p$-dimensional Brownian motion, independent of the $I_{1}(f^{(i)},0,t_{i})$’s, and $\widehat{\mathcal{C}}_{s}$ is the $p\times p$ covariance matrix of $\{I_{2}(f^{(i)},0,t_{i}),i\in[p]\}$, i.e. $\widehat{\mathcal{C}}_{ij}$ is defined analogously to the first line of (5.21). Note that $\widehat{\mathcal{C}}_{s}$ is random. We now show that we can approximate $I_{2}(f^{(i)},0,t_{i})$ with a Gaussian process by replacing $\widehat{\mathcal{C}}_{s}$ with its deterministic approximation. Define
@@
and the $p$-dimensional Gaussian process
$\big(\textbf{Z}_{s}\big)_{i}:=\int_{0}^{t_{i}}\big(\mathcal{C}_{s}^{1/2}\,{\mathrm{d}}\mathfrak{b}_{s}\big)_{i},\qquad\quad i\in[p].$ (5.24)
By analogous computations to (5.21) it follows that $\widehat{\mathcal{C}}_{ij}=\mathcal{C}_{ij}+\mathcal{O}(N^{-c})$, with very high probability. To estimate the difference between $(\textbf{Z}_{s})_{i}$ and $I_{2}(f^{(i)},0,t_{i})$ we thus compute the quadratic covariation
$\begin{split}\left|\langle(\textbf{Z}_{s})_{i}-I_{2}(f^{(i)},0,t_{i}),(\textbf{Z}_{s})_{j}-I_{2}(f^{(j)},0,t_{j})\rangle_{t}\right|&=\left|\int_{0}^{t_{i}\wedge t_{j}}\big(\big[\widehat{\mathcal{C}_{s}}^{1/2}-\mathcal{C}_{s}^{1/2}\big]^{2}\big)_{ij}\,{\mathrm{d}}s\right|\\ &\leqslant\int_{0}^{t_{i}\wedge t_{j}}\big(\big[\widehat{\mathcal{C}_{s}}^{1/2}-\mathcal{C}_{s}^{1/2}\big]^{2}\big)_{ii}^{1/2}\big(\big[\widehat{\mathcal{C}_{s}}^{1/2}-\mathcal{C}_{s}^{1/2}\big]^{2}\big)_{jj}^{1/2}\,{\mathrm{d}}s\\ &\leqslant\int_{0}^{t_{i}\wedge t_{j}}\mathrm{Tr}\big|\widehat{\mathcal{C}_{s}}-\mathcal{C}_{s}\big|\,{\mathrm{d}}s\\ &\leqslant\int_{0}^{t_{i}\wedge t_{j}}\sum_{i,j=1}^{p}\big(\widehat{\mathcal{C}_{s}}-\mathcal{C}_{s}\big)_{ij}\,{\mathrm{d}}s\lesssim N^{-c},\end{split}$ (5.25)
for a small $c>0$, where in the second inequality we used that $\mathrm{Tr}\big[(\widehat{\mathcal{C}_{s}}^{1/2}-\mathcal{C}_{s}^{1/2})^{2}\big]\leqslant\mathrm{Tr}\big|\widehat{\mathcal{C}_{s}}-\mathcal{C}_{s}\big|$ (see e.g. [Pos70, Section 3]). Then, from the BDG inequality it follows that $I_{2}(f^{(i)},0,t_{i})=(\mathbf{Z}_{s})_{i}+\mathcal{O}(N^{-c})$ with very high probability for all $i\in[p]$. In particular, this shows that, modulo a negligible error $N^{-c}$, the $I_{2}(f^{(i)},0,t_{i})$ are jointly a multivariate Gaussian random variable.
Now a Wick theorem as in (3.9) for the product (5.22) easily follows. In fact, (5.24)–(5.25) readily imply a Wick theorem for $I_{2}(f^{(i)},0,t_{i})$; we are thus left only with products of the $I_{1}(f^{(i)},0,t_{i})$’s . Then, by [CipErdSch2023, Proposition 3.3], we also conclude that the $I_{1}(f^{(i)},0,t_{i})$’s satisfy a Wick theorem as they consists of product of resolvents evaluated at the same time. This concludes the proof of of Proposition 3.3.
@@
$\begin{split}&\phantom{{}={}}\int_{t_{1}}^{t^{\prime}}\mathbb{E}\left[L_{N}(F,s_{2})\sum_{k}(-\frac{1}{2}\nabla G(\sigma_{k}(t))\cdot\sigma_{k}(t)+\frac{g(\sigma_{k}(t))}{4}\,\frac{\mathscr{O}_{kk}(t)}{N})\right]{\rm d}t\\ &=-\int_{t_{1}}^{t^{\prime}}\frac{1}{16\pi^{2}}\int_{\mathbb{C}^{2}}f_{v,a}(z)g_{v,a}(w)\frac{c_{v}}{c_{v}|t-s_{2}|+|z-w|^{2}}\,{\mathrm{d}}z{\mathrm{d}}w{\rm d}t+\operatorname{O}(N^{-4a-c+{\varepsilon}})\end{split}$ (5.29)
@@
which together with 5.29 gives
@@ Appendix A Stieltjes transform
$|m(z)|\leqslant\frac{1}{{\rm dist}(z,\operatorname{supp}(\mu))},\quad\quad|\partial_{z}m(z)|\leqslant\frac{{\rm Im}[m(z)]}{{\rm Im}[z]},\quad|\partial_{z}^{p}m(z)|\leqslant\frac{p!{\rm Im}[m(z)]}{{\rm dist}(z,\operatorname{supp}(\mu))^{p-1}{\rm Im}[z]}.$
@@
$|\partial_{z}m(z)|=\left|\int_{\mathbb{R}}\frac{{\rm d}\mu(x)}{(x-z)^{2}}\right|\leqslant\int_{\mathbb{R}}\frac{{\rm d}\mu(x)}{|z-x|^{2}}=\frac{1}{{\rm Im}[z]}\int_{\mathbb{R}}\frac{{\rm Im}[z]{\rm d}\mu(x)}{|z-x|^{2}}=\frac{{\rm Im}[m(z)]}{{\rm Im}[z]}.$
@@
$|\partial_{z}^{p}m(z)|=\left|\int_{\mathbb{R}}\frac{p!{\rm d}\mu(x)}{(x-z)^{p+1}}\right|\leqslant\int_{\mathbb{R}}\frac{p!{\rm d}\mu(x)}{|z-x|^{p+1}}\leqslant\int_{\mathbb{R}}\frac{p!{\rm d}\mu(x)}{|z-x|^{2}{\rm dist}(z,\operatorname{supp}(\mu))^{p-1}}=\frac{p!{\rm Im}[m(z)]}{{\rm dist}(z,\operatorname{supp}(\mu))^{p-1}{\rm Im}[z]},$
@@
$|\partial_{z}m(z)|\lesssim\max|\partial_{z}\widetilde{m}(u)|+\frac{\sqrt{\phi{\rm Im}[m(z)]}}{\eta}.$ (A.2)
@@
$|\partial_{u}{\rm Im}[m(u)]|\leqslant|\partial_{u}m(u)|\leqslant|\frac{{\rm Im}[m(u)]}{{\rm Im}[u]},$
@@
$|m(w)-m(z)-(w-z)\partial_{z}m(z)|\leqslant\frac{1}{2}|w-z|^{2}\sup_{u}|\partial_{u}^{2}m(u)|$
@@
$|\partial_{z}m(z)|$ $\leqslant\frac{|m(w)-m(z)|}{|w-z|}+\frac{1}{2}|w-z|\sup_{u}|\partial_{u}^{2}m(u)|\lesssim\max|\partial_{u}\widetilde{m}(u)|+\frac{\phi}{|w-z|}+|w-z|\frac{{\rm Im}[m(z)]}{{\rm Im}[z]^{2}}$
$\lesssim\max|\partial_{u}\widetilde{m}(u)|+\frac{\sqrt{\phi{\rm Im}[m(z)]}}{{\rm Im}[z]},$
@@ Appendix B Well-posedness
Still following [RogShi1993], consider $\phi({\bf{x}})=-\sum^{*}_{i,j}\log|x_{i}-x_{j}|$, where here and in the following $\sum_{i,j}^{*}$ (resp. $\sum^{*}_{i}$) means summation over couples $N\leqslant i\neq j\leqslant N$ such that $i,j\neq 0$ (resp. $N\leqslant i\leqslant N$ such that $i\neq 0$). Let ${\bf{y}}_{t}:={\bf{x}}_{t\wedge\tau_{\varepsilon}}$, then Itô’s formula gives
@@
so that the above parenthesis is smaller than ($\sum^{*}_{i,j,k}$ below means summation over all $N\leqslant i,j,k\leqslant N$ such that $i,j,k$ are all distinct and non-zero)
$\sum_{i}^{*}\left(\sum^{*}_{j:j\neq i}\frac{1}{y_{i}-y_{j}}\right)^{2}+\sum^{*}_{i,j}\frac{1}{(y_{i}-y_{j})^{2}}=-\sum^{*}_{i,j,k}\frac{1}{(y_{i}-y_{j})(y_{i}-y_{k})}=0,$